OpenStudy (anonymous):

if the volume V=4/3 pi r^3 of a sphere is changing at a rate of 40 cubic cm per min, find the rate at which the surface are A=4pi r^2 is changing, when V=400/3 pi cubic cm

OpenStudy (anonymous):

\[r=\sqrt[3]{3V/4 \pi}=\sqrt[3]{3(400 \pi/3)/4 \pi}=\sqrt[3]{100}\] A=3V/r dA/dt=d(3V/r)/dt \[dA/dt=d(3V/\sqrt[3]{100})/dt=(3/\sqrt[3]{100})(dV/dt)=3/\sqrt[3]{100}(40cm^3/\min)\]

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

how did you get A=3V/r?